OpenStudy (anonymous):

If g(x) = .5x + 3, find the value of g(0) + 2. A. 5 B. 4 C. 5.5 D. 7

OpenStudy (anonymous):

g(0) = .5 (0) + 3 = 3 g(0) + 2 = 3 + 2 = 5

OpenStudy (anonymous):

What is the slope of the line passing through (0,3) and (4,5)? A. 0.5 B. 2 C. -2 D. -0.5

OpenStudy (anonymous):

Cna u answer this for me too?

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Skylark joined a karate class. She paid $30 to join plus $25 per lesson. What is the rate in this problem? A. total cost B. original payment of $30 C. $25 per lesson

OpenStudy (anonymous):

What about this one

OpenStudy (anonymous):

The rate (gradient) is C

OpenStudy (mertsj):

The rate is $25 per lesson

OpenStudy (anonymous):

What is the correlation between the two quantities? The energy bill for running an airconditioner and the temperature outside. A. Positive B. Negative C. No correlation

OpenStudy (anonymous):

wat about this one?

OpenStudy (anonymous):

The hotter it is outside. The more likely you are to use an air conditioner. Therefore if one thing increases whilst the other increases. The correlation would be positive.

OpenStudy (anonymous):

Under water, pressure increases 4.3 pounds per square inch (psi) for every 10 feet you descend. This is represented by the equation p = 0.43d + 14.7, where p is the psi and d is the depth in ft. Find the pressure at 200 feet. A. 215.13 psi B. 860 psi C. 100.7 psi

OpenStudy (mertsj):

Can you replace d with = 200?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

That should give you p

OpenStudy (anonymous):

okay is the answer be C?

OpenStudy (mertsj):

That's what I got. Good for you. (Note: no exclamation points)

OpenStudy (anonymous):

lol i see tht. thnk u!! :)

OpenStudy (anonymous):

Determine if the following is always, sometimes, or never true. y = -1/5 x + 3 is a function. A. Always B. Sometimes C. Never

OpenStudy (mertsj):

A function means that you never get two ordered pairs with the same x value. For example you could not have (3,7) and (5,7) in a function.

OpenStudy (mertsj):

So you look at you equation and think: Will any x value give me two y values? or will each x give me only 1 y value?

OpenStudy (mertsj):

And another way to think of it is to think of its graph. Do you know that the graph of that equation will be a line?

OpenStudy (anonymous):

So will it be B?

OpenStudy (mertsj):

No. It is a line so it passes the vertical line test and will always be a function.

OpenStudy (anonymous):

Oh idk how to do these very well.

OpenStudy (mertsj):

Do you know that the graph of that equation will be a line?

OpenStudy (anonymous):

....Um.... no.

OpenStudy (mertsj):

y=mx+b is a line. Just like the other question you just posted. m is the slope and b is the y intercept.

OpenStudy (anonymous):

oh okay